OpenStudy (maeganr):

is - x + y = 3 a direct variation?

OpenStudy (anonymous):

Can it be rewritten in the form y = kx? If not, then it is not direct variation. Hint: can you solve your equation for y? Once you do, is there a constant you add to it?

OpenStudy (maeganr):

ok

OpenStudy (maeganr):

It can be rewritten as that, therefore it is not right?

OpenStudy (anonymous):

what does your rewritten equation look like?

OpenStudy (maeganr):

y = 3X

OpenStudy (anonymous):

Be careful, remember you can only combine like terms. If you solve this equation for y you'll get: -x + y = 3 +x +x add x to both sides to get y by itself y = 3 + x 3 and x are not "like terms" -- they have different variables/powers y = x + 3 rewrite to put x value first You were really close the first time and I like that you knew it was direct variation because of that. However, this is not because we're adding a constant (3) to x. In order for this to be direct variation we would just need a number in front of x without adding anything on.

OpenStudy (maeganr):

Oh ok, I thank you.